INFRARED SPECTROSCOPY OF THE PROPARGYL RADIAL IN SUPERFLUID LIQUID HELIUM DROPLETS

\underline{JEREMY MERRITT}, JOCHEN KÜPPER, and ROGER E. MILLER, University of North Carolina, Department of Chemistry, Chapel Hill, NC 27599.

An effusive pyrolysis source for the generation of a continuous beam of radicals suitable for the helium droplet pick-up process is presented as applied to the pyrolysis of Propargyl-Bromide to produce the Propargyl Radical. Rotationally resolved spectra are reported for the ν_{1} vibrational mode of the Propargyl radical in helium droplets. Stark spectra were also recorded that allowed for the first experimental determination of the permanent electric dipole moment of Propargyl, namely 0.15 D and 0.148 D for ground and excited state, respectively, in good agreement with previously reported ab initio results of 0.14 D ^{a}. The infrared spectrum of the ν_{1} mode of Propargyl-Bromide and future applications of these methods for the production of novel radical clusters will also be discussed.

aP. Borschwina, R. Oswald, J. Flügge, and M. Horn, Z. Phys. Chem. 188, 29 (1995).